OpenStudy (anonymous):

If -4 is a root of the quadratic equation x^2 + px - 4 = k and the quadratic equation x^2 + px + k = 0 has equal roots, find the value of k.

OpenStudy (phi):

and the quadratic equation x^2 + px + k = 0 has equal roots, are you sure there is no typo? as stated you get irrational numbers for k. it works better if this 2nd equation is x^2 +px -k =0

OpenStudy (anonymous):

that is how the question is given in the book howeverr the answer given is 9/4 !!

OpenStudy (phi):

OK, in that case, the typo is in the first equation (and makes the problem easier) they must mean If -4 is a root of the quadratic equation x^2 + px - 4 = 0 (not k) now you only have to find p, which can be done by substituting x=-4 and solving for p

OpenStudy (phi):

once you know p, you can "complete the square" to find what k is in the 2nd equation.

OpenStudy (anonymous):

@phi Thanks !! I'll try it out and see....